OpenStudy (anonymous):

i have this math problem im stuck they want to know how many hours she exercise the hours are 1/4, 1/2, 1 and 1/4, 0, 2 and 1/4 whats the answer

OpenStudy (anonymous):

1/4 hour= 1/4x60 = 15 minutes. 1/2 hour= 1/2x60 = 30 minutes. 1 hour = 1x60 minutes = 60 minutes. 1/4 hour = 15 minutes. 0 hour= 0 minutes. 2hours= 120 minutes. 1/4 hours= 15 minutes. Adding them all in minutes: 255 minutes. To find the time in hours, just divide by 60 because 60 minutes gives us 1 hour. So, 255/60 hours = 17/4 hours. is the answer.

OpenStudy (radar):

Just add the hours: 3/4 + 1/2 +3 hours 3 5/4 4 1/4 hours